3.2 Suppose that the residents of Windhoek spend all of their income on cauliflower,

broccoli and carrots. In 2019 they buy 100 heads of cauliflower for N$2000, 50 bunches

of broccoli for N$750 and 500 carrots for N$500. In 2020 they buy 75 heads of

Cauliflower for N$2250, 80 bunches of broccoli for N$1200 and 500 carrots for N$1000.

Suppose that the base year is 2019.

(a) Find nominal GDP and real GDP for 2019 and 2020. (4)

(b) What is the CPI in both years? (6)

(c) What is the inflation rate? ( 4)

(d) Find real growth rate (2)

(e) Which do you think has a greater effect on the CPI: a 10 percent increase in the

price of chicken or a 10 percent increase in the price of cars? Why? (4)


1
Expert's answer
2021-05-19T20:24:08-0400

a) The year 2019 and 2020 are the economic years for which the nominal and real GDP are calculated. The nominal and base year would be similar for 2019 as 2019 is a base year.

"Nominal \space GDP_{2019}= Real \space GDP_{2019}=\$2,000+\$750+\$500\\ =\$3,250"

The real GDP is the GDP which is calculated at the base year prices. The prices for one unit of broccoli, carrots and cauliflower in 2019 would be $20, $15 and $1 respectively.

"Nominal\space GDP_{2020}=\$2,250+\$1,200+\$1,000\\=\$4,450"

"Real \space GDP_{2020}=\$20×75+\$15×80+\$1×500\\=\$1,500+\$1,200+\$500\\=\$3,200"

b)

The CPI measures the cost of the market basket in the current year and base year and is the ratio of the cost of market basket in current year to the base year times 100.


"CPI_{2019}=\frac{\$2,000+\$750+\$500}{\$2,000+\$750+\$500}×100\\=1×100\\=100"


"CPI_{2020}=\frac{\$2,250+\$1,200+\$1,000}{\$20×75+\$15×80+\$1×500}\times 100"


"=\frac{\$4,450}{\$1,500+\$1200+\$500}\times 100"


"=\frac{\$4,450}{\$3,200}\times100"


"=139.06"


c)

Thus inflation rate is the percentage change in price level and the CPI being a measure of inflation the inflation would be equal to the percentage change in CPI from 2019 to 2020.

"Inflation\space Rate=\frac{CPI_{2020}−CPI_{2019}} {CPI_{2019}}×100"


"=\frac{139.06−100}{100}×100"


"39.06\%"


(d)

"Growth\space rate= \frac{GDP \space in \space 2020- GDP\space \space in \space 2019}{GDP\space in\space 2019}"


"=\frac{4450-3250}{3250}\\=0.3692\times 100\\=36.92\%"


Therefore the growth rate is 36.92%


(e)

Since chicken is a larger component of the average consumer's market basket, a 10% rise in the price of chicken has a greater impact on the consumer price index than a 10% increase in the price of a vehicle and car is not part of the consumer basket, so the increase in the price of the car will not show much effect on the CPI.